The time t required to drive a fixed distance varies inversely as the speed r. It takes 5 hrs at 55 mph to drive a fixed distance. How long would it take at 40 mph?
Answer by nyc_function(2645)

t varies inversely as r is written as t = k/r where k is the constant of proportionality.

We first need to find k plugging 5 for t and 55 for r.
5 = k/55
5*55 = k
275 = k
We now need to find t when k = 275 and r = 40.
t = k/r
t = 275/40
t = 6.875 hours